Pleasantville Hospital is a 100-bed hospital. The data table shows A&C data for November. A&C Data...

  1. Pleasantville Hospital is a 100-bed hospital. The data table shows A&C data for November.

A&C Data for November

Admissions

675

Discharges

683

Inpatient Service Days

2,653

Discharge Days

2,685

Calculate the following:

  1. Average length of stay for A&C in November

  1. A&C occupancy percentage for November

Please show your steps
Answer #1

a.

Average length of stay is the index which indicates the average period in hospital per patient admitted.

A=Total number of bed days in the year(H) divided by the number of admission in the same year(A)

=2653/675

=3.9

=4days.

b

Occupancy percentage=Total inpatient service days in a period/Total bed count days

Total bed count days=Bed count ×no of days in a period

Bed count=100

No of days=30 for November

Total inpatient service=2653

Total bed count days=100×30=3000

2653÷3000×100=88.4%